A full rotation is measured as 360°. A dial moves through a third of a rotation. How many degrees does it need to move backwards to reach a fifth of a rotation?


1
Expert's answer
2021-05-24T16:22:34-0400

At "3^{rd}" Rotation Dial moves "= 360\times 3=1080^{\circ}"


At "5^{th}" Rotation Dial moves "= 360\times 5=1800^{\circ}"


To move backwards to reach a fifth of a rotation Dial must moves-


"=1800-1080=720^{\circ}"